# Motion Primitives Planning For Center-Articulated Vehicles

### v1

@@ IV-B Real-world Experiments
We evaluated the effectiveness of our method on SAHA in an unstructured test field, setting with artificial obstacles. The environment is unknown beforehand, the lidar installed on the right front side of the vehicle (see Fig. 6) is used for state estimation and obstacle detection in real-time. In our experiments, the maximum speed $v_{max}$ of the vehicle is set to $1.0m/s$. In the first test, SAHA starts from the entrance of the field and heads to the goal near the edge of the field. On the way to the goal, there are columnar obstacles and block obstacles of different heights. To avoid these obstacles, SAHA’s final trajectory took on an ”S” shape, as shown in Fig. 1. We also test the planner’s ability to choose bi-directional trajectories to reach the goal set in the ”Unreachable Zone”, which is presented on the two sides of the vehicle due to its limited steering angles. As shown in Fig. 9, the vehicle successfully reaches the goal by first reversing and then moving forward.
V Conclusion
This work introduced a motion primitives-based planning strategy and a pose-stabilizing controller designed specifically for CAVs in unstructured environments. Our method involves a receding horizon planner based on offline-generated motion primitives tailored to the vehicle’s kinematic constraints, and a pose-stabilizing controller leveraging the kinematic model to improve path-tracking performance. Through rigorous simulation and real-world testing, including comparative analyses against the baseline cubic spline-based primitives and modified pure-pursuit control, as well as a real-world validation on SAHA, the method demonstrated significant improvement compared to previous methods and solid effectiveness in real-world autonomous navigation.

### v2

@@ V Conclusion
This work introduces a motion primitives-based planning strategy and a pose-stabilizing controller designed specifically for CAVs in unstructured environments. Our method involves a receding horizon planner based on offline-generated motion primitives tailored to the vehicle’s kinematic constraints, and a pose-stabilizing controller leveraging the kinematic model to improve path-tracking performance. We conduct rigorous simulation and real-world testing, including comparative analyses against baseline methods such as cubic spline-based primitives and modified pure-pursuit control, as well as real-world validation on SAHA. Through these evaluations, our method demonstrates significant improvements over previous methods and solid effectiveness in real-world autonomous navigation.